A perfect gas at 27 degree celcius is heated at constant pressure so as to double its volume The temp of gas will be

Grade:12th pass

1 Answers

Arun
6 years ago
T1 = 27 +273 =300 K 

V1 =V

V2=2V 

T2 =?

use  P1V1/T1  = P2V2/T2 

GIVEN P1 =P2

SO V1/T1  =V2/T2 

V/300 = 2V/T2

T2 =600K 

= 327 C